### v1

@@ 3 Agent-based Autonomous Supply Chain (A2SC)
Agents are goal-directed computation entities that act autonomously based on their designated roles, while also communicating and coordinating with other agents as needed [Wooldridge and Jennings, 1995, Fox et al., 2001]. A Multi-Agent System (MAS) consist of multiple such agents, interacting to address complex problems that exceed the capacities or knowledge of any individual agent. MAS approaches are, therefore, well-suited for addressing problems involving: 1) Multiple (semi)-autonomous entities; 2) Distributed knowledge among many entities; 3) The need for entities to coordinate to accomplish specific tasks. Supply chains, which consist of a set of distributed business entities, exemplify such problems. Consequently, the MAS approach is an ideal choice for constructing ASC systems by coordinating automated functions and processes. The resulting ASCs are referred to as agent-based autonomous supply chains (A2SC).
@@ 3.1 Multi-Agent System Approach
Various methodologies have been proposed to analyse and design multi-agent systems, and the Gaia methodology [Wooldridge et al., 2000, Zambonelli et al., 2003] has gained recognition for its simplicity and effectiveness. However, Gaia has certain limitations; it lacks completeness in abstraction and thus has no capacity to identify organisational rules and model environment. Consequently, Gaia could only be used for closed communities with cooperating agents. To address these limitations, Juan et al. [2002] introduced ROADMAP, an extension of Gaia that incorporates additional models into its framework. Additionally, Zambonelli et al. [2003] extended Gaia by using organisational abstractions, resulting in an extended Gaia capable of analysing and designing complex open systems with self-interested agents. In contrast to ROADMAP, which is considered preliminary, this extended Gaia methodology is comprehensive and provides coherent integration with the original Gaia approach. We therefore chose to employ Gaia-based methodologies for the analysis and design of ASC systems, customising and adapting them to the specific context of supply chains, which are open systems and often involve self-interested agents.
Figure 1 illustrates the adapted methodology for the analysis and design of ASC systems, which is derived from the Gaia and ROADMAP methodologies. This methodology comprises five key stages, each of which involves specific primary tasks and output models, as outlined below:
@@
    Implementation Design: While the previous stages mainly focus on analysing and designing the system’s structures without depending on specific implementation details, this stage is dedicated to designing elements that facilitate the system’s dynamics, particularly the flow of information through messaging. It involves three critical aspects related to agent interaction: knowledge representation, interaction protocols, and communication languages. Although these aspects do have relevance to implementation, they are included in this model as an independent design stage because the coherence of a MAS relies heavily on effective messaging. Other concrete system implementation aspects are excluded from this stage.
The methodology outlined above describes the two main phases that come before the development phase of implementing an ASC system. Each of these phrases comprises several stages, with each stage encompassing a set of key artefacts, as highlighted in blue in Figure 1. These artefacts can be broadly categorised into three groups based on their purpose: structural, social, and semantic. Structural artefacts, represented in light blue in Figure 1, are those models related to the composition and organisational aspects of a MAS. Social artefacts, highlighted in light green, pertain to the interactions and relationships among agents. Semantic artefacts, indicated in light violet, concern the understanding of the interaction content. These artefacts collectively contribute to the analysis and design of an ASC system, providing conceptual models for subsequent development phrase. In the following paragraphs, we delve into these artefacts by reorganising them into the four concrete issues when applying a MAS approach [Smith and Davis, 1981, Genesereth, 1997, Fox et al., 2001].
@@
In a MAS, as described in [Wooldridge and Jennings, 1995], multiple agents collaborate to achieve collective behaviours. These individual agents typically lack sufficient information or capability to independently solve the entire problem. Therefore, effective collaboration among these agents is essential to ensure the system operates coherently. To facilitate this collaboration, it is crucial to thoroughly decompose the problem and identify specific roles that agents should play in addressing the problem. Agents can be assigned one or more specific roles, defining their expected behaviours (or services) within the system. Additionally, it is important to organise these agents into a specific organisational structure that defines the topology of the interaction patterns and the control regime for activities within this organisation. The structural artefacts, which includes roles, agents and their services, and organisation as illustrated in Figure 1, are the main models for representing the composition of a MAS. While this issue tackles the static structural aspects of a MAS, the remaining three issues focus on its dynamics, including interactions and elements that facilitate communication.
@@
While the first two issues define the structure of a MAS, this issue deals with dynamics, specifically the flow of messages along the communication channels between its constituent agents. This includes establishing protocols that regulate interaction process and defining communication languages for representing messages. Some representative works in these areas include the Contract-Net protocol [Davis and Smith, 1983], and the English auction protocol [FIPA, 2001], and KQML [Finin et al., 1994], and FIPA-ACL [FIPA, 2002, Poslad, 2007].
@@
With the first three issues addressed, the flow of messages is facilitated; however, message comprehension remains unaddressed. To ensure that agents understand the content of messages, they must use a standardised vocabulary with teams that consistently refer to objects, functions, and relations relevant to a particular application. This consistent vocabulary, along with a commonly understood knowledge representation, constitute the fourth issue that requires attention when applying a MAS approach. A vocabulary may encompass multiple ontologies used to describe the shared body of knowledge within a specific domain.
These four issues are essential when adopting an MAS approach. The first two issues pertain to the structural aspects (composition and organisation) of an MAS, whereas the last two relate to its dynamics, specifically, interaction and communication. Commonly communication languages and interaction protocols allow agents to freely join or leave the MAS, enabling an open, dynamical computational environment. These four issues are a concrete refinement of the methodology illustrated in Figure 1. Despite the differences between supply chains and traditional MAS scenarios, these four issues provide valuable guidance for the analysis and design of a A2SC system.
3.2 Analysis and Design of A2SC
An ASC, as described in [Xu et al., 2022], is an self-operating supply chain comprised of a set of structural entities. Each of these entities may represent individual business entities in a supply chain and can manage multiple automated processes. These structural entities are crucial components in the formation of an ASC and are responsible for managing the three flows within the supply chain. Efficient operation of an ASC requires effective coordination among these structural entities. However, their coordination is challenging due to their inherently self-interested, distributed, heterogeneous, yet cooperative nature. Fortunately, MAS approaches are well-suited for building open systems that involve both cooperative and self-interested entities.
Given the diverse range of automated sub-systems involved in an ASC, often using technologies from various domains, delving into the specific designs of these systems is neither necessary nor feasible within the scope of this article. Instead, we treat these automated systems as readily available building blocks for the assembly of an ASC. We then primarily focus on the integration of these structural entities and their internal processes, enabling seamless and autonomous collaboration. In this section, we discusses the analysis and design of A2SC system using the MAS methodology presented in the previous section.
@@ 3.2 Analysis and Design of A2SC
In the context of A2SC, problem decomposition must take into account organisational boundaries. According to Xu et al. [2022], structural entities act as bridge points for business entities, allowing them to connect with external peers while managing their internal entities within a supply chain. Decomposing the supply chain into a set of structural entities, each represent multiple roles and potentially represented by an agent, aligns with the natural structure of the supply chain.
@@
Despite operating at a higher level of abstraction, this decomposition approach somewhat aligns with with existing literature. Previous studies, such as Swaminathan et al. [1998], Fox et al. [2001], Akanle and Zhang [2008], identified two main categories of agents: structural and functional, based on their organisational and functional roles. Structural agents represent the organisational entities involved in a supply chain, including stakeholders like suppliers, retailers, third-party logistics (3PL) providers, distribution centres, and manufacturing plants. On the other hand, functional agents are responsible for specific functionalities in a supply chain, such as supplier selection, order placement, and logistics arrangement. A structural entity represent an conceptual organisation, thus it representative agent may manage multiple these structural and functional agents.
@@
These decomposed agents collectively manage the operations of a supply chain through interactions. Agent interaction defines the potential interdependencies among agents in an MAS, thus influencing the MAS’s topology. Therefore, interactions among agents must be carefully designed to create a loosely-coupled, scalable MAS which agents focus more on computation rather than communication. In A2SCs, the management of the three flows — product, information, and financial — involves agents interacting with others to facilitate the inbound and outbound of movement of these flows. Developing appropriate interactions is crucial to ensuring the efficient coordination of these flows through a network of structural entity agents. During interactions, agents may exhibit two types of behaviours: proactive and reactive. Proactive behaviours involve agents initiating interaction to pursue their objectives, such as connecting other agents to request price data. Conversely, reactive behaviours are response to external events, such as handling incoming requests when receiving goods. Furthermore, interactions should be governed by protocols, which are the commonly recognised procedures or system of rules. These protocols manage various interaction processes, each regulating agents’ dialogue with others in different scenarios, such as supplier selection, procurement price negotiation, and advanced shipment notification.
@@
Agents interact through messages; and in an A2SC, the entire system functions through the exchange of these messages. For an A2SC to function efficiently, these messages must be fully comprehended both syntactically and semantically by all participating agents. Metaphorically, agents in an A2SC community must “speak” a common language to enable clear and unambiguous communication. This common language should exhibit consistency in three critical aspects:
@@
These three aspects collectively standardise the messages in terms of their structure, content representation, and semantics. To achieve universally understood messages, they should be explicitly specified using common languages. Information flow among agents in an A2SC, including quotations, purchase orders, delivery status, and invoices, relies on effective messaging. Consequently, careful considerations on interaction protocols, languages, and message representation is crucial for the successful implementation of an A2SC.
@@ 4 A Case Study: an Autonomous Meat Supply Chain
The previous section introduces the A2SC and outlines the methodology for its analysis and design. In this section, we present a detailed case study of A2SC implementation to illustrate the development of an ASC prototype using an MAS approach. We selected the meat supply chain — a particular perishable foods supply chain — as the case for study. Perishable foods, unlike other goods, are prone to spoilage due to inadequate storage conditions. To ensure the successful delivery of perishable foods, it is crucial to maintain proper storage conditions and minimise transit times. Monitoring the environmental conditions of trucks transporting perishable foods and swiftly addressing unforeseen events, such as traffic congestion, is of paramount importance.
In this prototype, we address the procurement of meat, specifically implementing an end-to-end automated meat procurement process. This procurement process involves several key stakeholders in the meat supply chain. However, it is important to not that our representation of meat supply chain is simplified and simulated (Figure 2). This simplified version excludes minor participants, such as secondary processors, which does not exactly mirror the complexity of real-world meat supply chain operations. Figure 2 provides an overview of the main structure of this meat supply chain. Our prototype primarily focuses on the central components highlighted in grey.
Furthermore, this prototype includes logistics and third-party logistics (3PL) since they are responsible for transportation, which are essential to PFSCs. For simplification purposes, we omitted farmers and end customers. The core procurement process involving suppliers, wholesalers, and retailers is sufficient for illustrating an initial A2SC implementation, and including additional entities would lead to unnecessary redundancy.
@@ 4.1 Scenario: the Cambridge Meat Company
From this scenario, we can further analyse the requirements for the prototype, aiming to develop an end-to-end automated process for meat procurement and wholesale. A crucial requirement for this prototype is to ensure the smooth movement of physical flow and its corresponding information flow. Given that modern financial transactions primarily rely on digital information, essentially constituting a specialised from of information flow, this prototype does not encompass financial flow considerations. Additionally, this prototype take into account inventory management, continuously updating it as goods are received and dispatched. The following sections first extract requirements and then detail the design, implementation, and showcase of the A2SC prototype.
@@ 4.2 Process Design
This prototype includes two primary processes: replenishment (the CMC procures meat from its suppliers to replenish its inventory) and wholesale (the retailer buys meat from the CMC). Figure 3 presents a three-dimensional diagram of the refined CMC scenario, with the two processes represented by the flow of blue and red arrows. In addition to the three main stakeholders (supplier, wholesaler, and retailer), logistics and 3PL, which provide logistics service, are included. Logistics service and inventory management, crucial for the flow of goods (meat), are considered in both processes. Both processes include logistics services, with sellers (supplier and wholesaler) responsible for delivery arrangements. Figure 4 illustrates the flowcharts of the these two processes, highlighting major steps. Different supply chain stakeholders are denoted by capital letters in brackets and are represented using different colours. The replenishment process (see Figure 3(a)) provides customers (wholesalers) with multiple delivery options, allowing them their preferred delivery service. However, the wholesale process (see Figure 3(b)) has a preselected delivery option by the wholesaler for their customers (retailers), making it less complex. Both processes involves multiple decision-makings points (denoted by diamonds), such as proposal evaluation, acceptance or rejection of proposals, and delivery option selection. Inventory-related functions, such as automatic inventory updates and replenishment, are integrated into both processes. Additionally, both processes incorporate logistics monitoring, which tracks real-time locations and ambient conditions of goods throughout delivery.

### v2

@@ 3.1 Agents and Multi-Agent System
Agents are goal-directed computation entities that act autonomously based on their designated roles, while also communicating and coordinating with other agents as needed (Wooldridge and Jennings, 1995; Fox et al., 2001). An MAS consist of a collection of such agents, interacting to address complex problems that exceed the capacities or knowledge of any individual agent. The MAS approach is therefore well-suited for addressing problems involving: 1) multiple (semi)-autonomous entities; 2) distributed knowledge among many entities; 3) the need for entities to coordinate to accomplish specific tasks. Supply chains, consisting of a set of distributed business entities, exemplify such problems. Consequently, the MAS approach is an ideal choice for constructing ASC systems by coordinating automated functions and processes. The resulting ASCs are referred to as A2SCs.
@@ 3.2 Multi-Agent System Approach
Various methodologies have been proposed to analyse and design multi-agent systems, and the Gaia methodology (Wooldridge et al., 2000; Zambonelli et al., 2003) has gained recognition for its simplicity and effectiveness. However, Gaia has certain limitations; it lacks completeness in abstraction and has no capacity to identify organisational rules and model environment. Consequently, Gaia could be only used for closed communities with cooperative agents. To address these limitations, Juan et al. (2002) introduced ROADMAP, an extension of Gaia that incorporates additional models into its framework. Additionally, Zambonelli et al. (2003) extended Gaia by using organisational abstractions, resulting in an extended Gaia approach capable of analysing and designing complex open systems with self-interested agents. In contrast to ROADMAP, which is considered preliminary, this extended Gaia is comprehensive and provides coherent integration with the original Gaia approach. We therefore chose to employ Gaia-based methodologies for the analysis and design of ASC systems, refining, customising, and adapting them to the specific context of supply chains, which are open systems and often involve self-interested agents.
The main adaptions includes: 1) we extract the models and elements from the Gaia methodology and its variants that are most relevant to the A2SC design and analysis; 2) we add implementation stage after the detailed design stage of the Gaia process, which has been specifically introduced in the Gaia methodology; and 3) we categorise the artefacts in the methodology into three groups: structural, social, and semantic, to better guiding the analysis and design of the static structure and dynamics of the A2SCs. Figure 1 illustrates the adapted methodology for analysing and designing A2SCs, with artefacts of different categories highlighted by blue, light green, and light violet, respectively. This adapted methodology comprises five key stages, one more stage than the Gaia methodology, each involving specific primary tasks and output models, as outlined below:
@@
    Implementation Design: While the previous stages mainly focus on analysing and designing the system’s structures without depending on specific implementation details, this stage is dedicated to designing elements that facilitate the system’s dynamics, particularly the flow of information through messaging. It involves three critical aspects related to agent interaction: knowledge representation, interaction protocols, and communication languages. Although these aspects do have relevance to implementation, they are included in this model as an independent design stage because the coherence of an MAS relies heavily on effective messaging. Other concrete system implementation aspects are excluded from this stage.
The methodology outlined above describes the two main phases (analysis and design phrases) that come before the development phase of implementing an ASC system. Each of these phrases comprises several stages, with each stage containing a set of key artefacts, as illustrated in Figure 1. Structural artefacts, represented in light blue in Figure 1, are those models related to the composition and organisational aspects of an MAS. Social artefacts, highlighted in light green, pertain to the interactions and relationships among agents. Semantic artefacts, indicated in light violet, concern the understanding of the interaction content. These artefacts collectively contribute to the analysis and design of an ASC system, providing conceptual guidelines for subsequent development phrase. In the following paragraphs, we delve into these artefacts by reorganising them into the four concrete issues when employing an MAS approach (Smith and Davis, 1981; Genesereth, 1997; Fox et al., 2001).
@@
In an MAS, as described in (Wooldridge and Jennings, 1995), multiple agents collaborate to achieve collective behaviours. These individual agents typically lack sufficient information or capability to independently solve the entire problem. Therefore, effective collaboration among these agents is essential to ensure the system operates coherently. To facilitate this collaboration, it is crucial to thoroughly decompose the problem and identify specific roles that agents should play in addressing the problem. Agents can be assigned one or more specific roles, defining their expected behaviours (or services) within the system. Additionally, it is important to organise these agents into a specific organisational structure that defines the topology of the interaction patterns and the control regime for activities within this organisation. The structural artefacts, which includes roles, agents and their services, and organisation as illustrated in Figure 1, are the main models for representing the composition of an MAS. While this issue tackles the static structural aspects of an MAS, the remaining three issues focus on its dynamics, including interactions and elements that facilitate communication.
@@
While the first two issues define the structure of an MAS, this issue deals with dynamics, specifically the flow of messages along the communication channels between its constituent agents. This includes establishing protocols that regulate interaction process and defining communication languages for representing messages. Some representative works in these areas include the Contract-Net protocol (Davis and Smith, 1983), and the English auction protocol (FIPA, 2001), and KQML (Finin et al., 1994), and FIPA-ACL (FIPA, 2002; Poslad, 2007).
@@
With the first three issues addressed, the flow of messages is facilitated; however, message comprehension remains unaddressed. To ensure that agents understand the content of messages, they must use a standardised vocabulary with teams that consistently refer to objects, functions, and relations relevant to a particular application. This consistent vocabulary, along with a commonly understood knowledge representation, constitute the fourth issue that requires attention when applying an MAS approach. A vocabulary may encompass multiple ontologies used to describe the shared body of knowledge within a specific domain.
These four issues are essential when adopting an MAS approach. The first two issues pertain to the structural aspects (composition and organisation) of an MAS, whereas the last two relate to its dynamics, specifically, interaction and communication. Commonly communication languages and interaction protocols allow agents to freely join or leave the MAS, enabling an open, dynamical computational environment. These four issues are a concrete refinement of the methodology illustrated in Figure 1. Despite the differences between supply chains and traditional MAS scenarios, these issues provide valuable guidance for the analysis and design of a A2SC system.
3.3 A2SC Analysis and Design
An ASC, as described in Xu et al. (2023b), is an self-operating supply chain comprised of a set of structural entities. Each of these entities may represent individual business entities in a supply chain and can manage multiple automated processes. These structural entities are crucial components in the formation of an ASC and are responsible for managing the three flows within the supply chain. Efficient operation of an ASC requires effective coordination among these structural entities. However, their coordination is challenging due to their inherently self-interested, distributed, heterogeneous, yet cooperative nature. Fortunately, the MAS approach is well-suited for building open systems that involve both cooperative and self-interested entities.
Considering the broad spectrum of automated processes involved in an ASC, often utilising technologies from various domains, delving into their specific designs of these processes is neither necessary nor feasible within the scope of this article. Instead, we assume these automated processes as readily available building blocks for the construction of an ASC. Our main focus then shifts to integrating these processes and their corresponding structural entities, enabling seamless and autonomous collaboration. This section discusses the concrete A2SC analysis and design using the MAS methodology presented in the previous section.
3.3.1 Decomposing ASC into Agents
@@
In the context of A2SC, problem decomposition must take into account organisational boundaries. According to Xu et al. (2023b), structural entities act as bridge points for business entities, allowing them to connect with external peers while managing their internal entities within a supply chain. Decomposing the supply chain into a set of structural entities, where each may represent multiple roles and potentially be represented by an agent, aligns with the natural structure of the supply chain.
@@
Despite operating at a higher level of abstraction, this decomposition approach is consistent to some extent with existing literature. Previous studies, such as those by Swaminathan et al. (1998); Fox et al. (2001); Akanle and Zhang (2008), identified two main categories of agents: structural and functional, based on their organisational and functional roles. Structural agents represent the organisational entities involved in a supply chain, including stakeholders such as suppliers, retailers, third-party logistics (3PL) providers, distribution centres, and manufacturing plants. Functional agents manage specific tasks in a supply chain, such as supplier selection, order placement, and logistics arrangement. A structural entity represent an conceptual organisation, thus it representative agent may oversee multiple structural and functional agents.
3.3.2 Designing Interactions and Messages
@@
These decomposed agents collectively manage the operations of a supply chain through interactions. Agent interaction defines the potential interdependencies among agents in an MAS, thus influencing the MAS’s topology. Therefore, interactions among agents must be carefully designed to create a loosely-coupled, scalable MAS which agents focus more on computation rather than communication. In A2SCs, the management of the three flows — product, information, and financial — involves agents interacting with others to facilitate the inbound and outbound of movement of these flows. Developing appropriate interactions is crucial to ensuring the efficient coordination of these flows through a network of structural entity agents. During interactions, agents may exhibit two types of behaviours: proactive and reactive. Proactive behaviours involve agents initiating interaction to pursue their objectives, such as connecting other agents to request price data. Conversely, reactive behaviours respond to external events, such as handling incoming requests when receiving goods. Furthermore, interactions should be governed by protocols, which are the commonly recognised procedures or system of rules. These protocols manage various interaction processes, each regulating agents’ dialogue with others in different scenarios, such as supplier selection, procurement price negotiation, and advanced shipment notification.
@@
Agents interact through messages; the entire A2SC system functions through the exchange of these messages. To function efficiently, these messages must be fully comprehended both syntactically and semantically by all participating agents. Metaphorically, agents in an A2SC community must “speak” a common language to enable clear and unambiguous communication. This common language should exhibit consistency in below critical aspects:
@@
These three aspects collectively standardise the messages in terms of their structure, content representation, and semantics. To achieve universally understood messages, they should be explicitly specified using common languages. Information flow among agents, including quotations, purchase orders, delivery status, and invoices, relies on effective messaging. Consequently, careful considerations on interaction protocols, languages, and message representation is crucial for the effective implementation of an A2SC.
@@ 4 A Case Study: an Autonomous Meat Supply Chain
The previous section introduces the A2SC and describes a methodology for its analysis and design. This section details a case study of A2SC implementation to illustrate the development of an ASC prototype using an MAS approach. We selected the meat supply chain — a particular perishable foods supply chain — as the case for study. Perishable foods, unlike other goods, are prone to spoilage due to inadequate storage conditions. It is crucial to maintain proper storage conditions and minimise transit times to ensure the successful delivery of perishable foods. Monitoring the environmental conditions of trucks transporting perishable foods and swiftly addressing unforeseen events, such as traffic congestion, is of paramount importance.
In this case, we focus on an integrated process — automated meat procurement — specifically implementing an autonomous process for procuring meat and transporting meat from producers to end customers. This procurement process involves several key stakeholders across the meat supply chain. However, it is important to note that our representation of meat supply chain is simplified and simulated (Figure 2). It excludes minor participants, such as secondary processors, which thus may not exactly mirror the complexity of real-world meat supply chain operations. Figure 2 presents an illustration of this supply chain. Our description mainly focuses on the central process highlighted in red.
Furthermore, this prototype includes logistics and 3PLs since they are responsible for transportation, which are essential to perishable food supply chains. The core procurement process involving suppliers, wholesalers, and retailers is sufficient for illustrating an initial A2SC implementation; and including additional entities would lead to unnecessary redundancy.
@@ 4.1 Scenario: the Cambridge Meat Company
From this scenario, we can further analyse the requirements for the prototype, aiming to develop an integrated automated process for meat procurement and wholesale. A crucial requirement for this prototype is to ensure the smooth movement of physical flow and its corresponding information flow. Given that modern financial transactions primarily rely on digital information, essentially constituting a specialised from of information flow, this prototype does not encompass financial flow considerations. Additionally, this prototype take into account inventory management, continuously updating it as goods are received and dispatched. The following sections first extract requirements and then detail the design, implementation, and showcase of the A2SC prototype.
Figure 3: Illustration of the CMC scenario.
@@ 4.2 Process Design
This prototype includes two primary processes: replenishment (the CMC procures meat from its suppliers to replenish its inventory) and wholesale (the retailer buys meat from the CMC). Figure 3 presents a diagram of the refined CMC scenario, with the two processes highlighted by the flow of blue and red arrows. In addition to the three main stakeholders (supplier, wholesaler, and retailer), logistics and 3PL, which provide logistics service, are included. Logistics service and inventory management, crucial for the flow of goods (meat), are considered in both processes. Both processes include logistics services, with sellers (supplier and wholesaler) responsible for delivery arrangements. Figure 4 illustrates the flowcharts of these two processes, highlighting major steps. Different supply chain stakeholders are denoted by capital letters in brackets and are represented using different colours. The replenishment process (see Figure 4(a)) provides customers (wholesalers) with multiple delivery options, allowing them their preferred delivery service. However, the wholesale process (see Figure 4(b)) uses a predetermined delivery option selected by the wholesaler for their customers (retailers), making it less complex. Both processes involves multiple decision-makings points (denoted by diamonds), such as proposal evaluation, acceptance or rejection of proposals, and delivery option selection. Inventory-related functions, such as automatic inventory updates and replenishment, are integrated into both processes. Additionally, both processes incorporate logistics monitoring, which tracks real-time locations and ambient conditions of goods throughout delivery.